a+b c c b c+a b a a b+c this is determinate problem i need easy solution in some operations

a+b c c


b c+a b


a a b+c


this is determinate problem i need easy solution in some operations

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
\begin{vmatrix} a+b & c &c \\ b& c+a & b\\ a& a & b+c \end{vmatrix}
r1=r1-(r2+r3)
\begin{vmatrix} 0 & -2a &-2b \\ b& c+a & b\\ a& a & b+c \end{vmatrix}
now by expanding the matrix we get
=2a(b2+bc-ab)-2b(ab-ac-a2)
=4abc